# Continuous Monitoring via Repeated Significance

### v1

@@ 9             Remove all zero-valued entries from $L$
for a value $10<w<$ 1 specified as part of the test plan.
IV Geometric $\alpha$-Spending
Figures 2 to 4 show $Z$ scores required for significance at each decision point for an unbounded test that allows stopping at the first decision point $t$ with a fraction at least $u$ of the decision points 1 to $t$ having $p$ values low enough for significance at their decision points, with probability of type 1 error at most $\alpha=0.05$. We plot $Z$ scores rather than $p$ values for ease of presentation, since $Z$ scores are familiar to many practitioners and have a 1-1 correspondence with $p$ values. The plotted $Z$ scores are the number of standard deviations from the mean of a standard normal such that the left and right tails together have probability $p$. (For comparison, a test with a single decision point that spends all of $\alpha=0.05$ on that decision point would require $Z\approx 1.96$ for $p=0.05$.)
The figures show that decreasing the “withdrawal” rate increases the number of decision points $t$ before reaching the minimum $Z$ score required for significance and decreases the rate of increase after reaching the minimum. Decreasing $w$ maintains more of the $\alpha$ budget for later decision points.
Together, the figures show that requiring more repetition makes requirements for significance at each decision point less stringent: as $u$ increases from $0.05$ to $0.10$ then to $0.5$ in Figures 2 to 4, the curves drop to lower $Z$ scores. (The exception is the first few values of $t$ on the left, since they require the same number of repetitions: $\lceil ut\rceil=1$ until $t>\frac{1}{u}$.)
Figure 2: For $u=0.05$, $\mathbf{Z}$ score for significance at each decision point. Requiring 5% of decision points to have significant results before ending the test makes the minimum required $Z$ scores about 3.3. A smaller “withdrawal” rate $w$ increases the number of decision points $t$ before reaching the minimum.
Figure 3: For $u=0.1$, $\mathbf{Z}$ score for significance at each decision point. Requiring 10% of decision points to have significant results before ending the test makes the minimum required $Z$ scores about 3.1 – lowering the curves from Figure LABEL:fiz_zgeo005 for $u=0.05$ by requiring more repetition.
Figure 4: For $u=0.5$, $\mathbf{Z}$ score for significance at each decision point. Requiring 50% of decision points to have significant results before ending the test makes the minimum required $Z$ scores about 2.6, compared to about 3.1 for $u=0.1$ and 3.3 for $u=0.05$ (as shown in Figures 3 and 2).
Figures 5 and 6 show results for larger numbers of decision points, both with $u=0.1$, as in Figure 3 (all with $\alpha=0.05$). To be clear, all those figures show results for unbounded tests, so they all have unlimited numbers of decision points, what is different is the number of initial decision points shown in each plot: ten thousand, a million, then a billion. Note that, after adjusting the values of $w$ in inverse proportion to the number of decision points plotted, the resulting curves are very similar, including their minimum $Z$ scores. So, by selecting $w$ appropriately, we can hold $u$ constant and have nearly the same curve, but stretched horizontally, with the same test effectiveness in terms of the $p$-values required for significant results. That said, increasing $t$ while holding $u$ constant maintains the same repetition requirement in terms of fraction of decision points, but increases it in absolute terms.
Figure 5: For $u=0.1$ and over the first million decision points, $\mathbf{Z}$ scores for significance. The “withdrawal” rates $w$ in this figure are one hundredth those in Figure 3, which is over the first ten thousand decision points – one hundredth of a million. The result is similar curves, with similar minimum $Z$ scores required for significance.
Figure 6: For $u=0.1$ and over the first billion decision points, $\mathbf{Z}$ scores for significance. The “withdrawal” rates $w$ in this figure are one thousandth those in Figure 5, which is over the million decision points, giving similar $Z$ score curves.
Figure 7 compares using geometric $\alpha$ spending with repetition requirement $u=0.2$ to an always-valid method based on autocorrelation among averages [14] (their Equation 8, page 8). The $Z$ score required for significance for that method is
@@ IV Geometric $\alpha$-Spending
with $\rho$ a parameter that controls the $t$ value at which the $Z$ score reaches a minimum, like $w$ for our method. Both methods apply to unbounded tests.
@@
Figure 7: Comparison to another always-valid method. For $u=0.2$, requiring repetition gives less stringent requirements for significance than a method based on autocorrelation for averages [14].
V Nearly Impossible is Possible
Figures 2 to 6 all have $Z$ score curves that increase after reaching a minimum, prompting the question: “Is there a way to spend $\alpha$ that gives a flat curve for an unbounded test, keeping the required $p$-value from shrinking to zero, or, equivalently, the required $Z$ score from becoming unbounded as $t\rightarrow\infty$?”
@@ V Nearly Impossible is Possible
Now we show that the answer to the question is a negative of the form: “It is not quite possible, but nearly so.”
@@
$=\frac{c}{\alpha u}\left(\frac{1}{t+\frac{1}{u}}\right).$ (15)
@@
diverges for $0\leq v\leq 1$ and converges for $v>1$. So in Expression 18, the first sum diverges and the second converges, implying that the expression diverges. ∎
In short, we cannot have $x_{t}$ proportional to $\frac{1}{t}$ because its $p$-series diverges. However, we can use $x_{t}$ proportional to any $\frac{1}{t^{v}}$ for any $v>1$, since its $p$-series converges. So, in a sense, Theorem 2 is the most nearly possible impossible result possible.
@@
Figures 8 shows the $Z$ scores required for significance, using $p$-series $\alpha$ spending with different exponents $v$, with
$x_{t}=\frac{1}{\zeta(v)t^{v}},$ (25)
so
$\delta_{t}=\frac{\alpha\lceil ut\rceil}{\zeta(v)t^{v}}.$ (26)
As $v\rightarrow 1$, the curve of required $Z$ scores becomes flatter. Figure 9 shows the same results in detail for the first hundred decision points. Note that as $v\rightarrow 1$, the $Z$ scores decrease for the initial few decision points, and the curves are not very flat initially.
Figure 8: For $p$-series $\alpha$ spending, $\mathbf{Z}$ scores for significance. Using $p$-series spending, the curves flatten as $v$ approaches $1$ from the left, though we cannot set $v$ to exactly one without overspending $\alpha$, by Theorem 2. ($u=0.1$ and $\alpha=0.05$.)
Figure 9: Detail: for $p$-series $\alpha$ spending, $\mathbf{Z}$ scores for significance. This is the data from Figure 8, for the first hundred decision points. The curves are not flat for the first few decision points, so $\alpha$ is, in a sense, “overspent” initially.
@@
$x_{t}=\frac{1}{\left(\zeta(v)-h(s,v)\right)\left(t+s\right)^{v}},$ (28)
and so
$\delta_{t}=\frac{1}{\left(\zeta(v)-h(s,v)\right)\left(t+s\right)^{v}}\alpha\lceil ut\rceil.$ (29)
Figures 10 shows the results of removing some initial terms from the $p$-series to produce $x_{t}$. The curve for $s=0$ has no terms removed; it is the curve for $v=1.1$ from Figure 8. Note that removing more terms lowers the required $Z$ scores for significance for most of the $t$ values. Figure 11 shows the same data, but just for the first hundred decision points, where the ordering is reversed: removing more terms increases required $Z$ scores for the initial decision points.
@@
Figure 11: Detail: headless $p$-series $\alpha$ spending: $\mathbf{Z}$ scores for significance. Plotting the data of Figure 10 shows that the lower $Z$ scores on the right of that plot are at the expense of raised $Z$ score requirements for the first few decision points.
VI Discussion
We have explored different ways to spend $\alpha$ with repetition requirements for unbounded tests, showing that it is possible to adjust the levels of significance required at each decision point by requiring more repetition. We have also shown that it is impossible to have a significance requirement that does not eventually become more stringent for an unbounded test, but we can stave that off for an arbitrary number of decision points by setting parameters appropriately.
An earlier paper [16] that focused on bounded tests discussed how to balance control of type 1 and type 2 error, by spending a portion of $\alpha$ on a single loose requirement for significance at the end of the test, which is reminiscent of traditional sequential testing strategies, such as the Haybittle-Peto boundary [28, 2]. It is a challenge to imagine something similar for unbounded tests, since their ending decision points are not known a priori. Perhaps spending a portion of $\alpha$ for loose requirements at a (possibly unbounded) subsequence of the decision points, either with or without a repetition requirement, would prove useful.
It would also be interesting to explore ways to combine repetition requirements with other approaches to continuous monitoring and always-valid bounds [10, 11, 12, 13, 14]. In some cases, it may be possible to improve both types of bounds by combining their strategies. The challenge is that non-repetition methods sometimes make use of side conditions that are proven empirically in some sense with repetition requirements – we need to analyze how these approaches overlap to understand how they can complement each other beyond mere duplication, if possible.

### v2

@@ 4 Geometric $\alpha$-Spending for Unbounded Tests
for a value $0<w<1$ specified as part of the test plan.
To find the maximum required $p_{t}$, take the derivative of the RHS of Inequality 26 with respect to $t$, set it to zero, and solve for $t$:
$t_{\min}\approx\frac{-1}{\ln(1-w)}.$ (27)
(We have $\approx$ rather than == only because we use $ut$ in place of $\lceil ut\rceil$ to find $t_{\min}$.)
Conversely, solve this for $w$ to find a value that makes the least stringent $p_{t}$ requirement occur at decision point $t$:
$w\approx 1-e^{-1/t}.$ (28)
This can be useful if we have some idea how long the test is likely to run before achieving statistically significant results. Equations 27 and 28 show that decreasing the “withdrawal" rate $w$ increases the number of decision points $t$ before reaching the maximum (least stringent) required $p_{t}$.
Figure 2 compares using geometric $\alpha$ spending with repetition requirement $u=0.2$ to an always-valid method based on autocorrelation among averages waudbysmith24 (their Equation 8, page 8), with probability of Type I error at most $\alpha=0.05$. (We plot $Z$ scores rather than $p$ values for ease of presentation, since $Z$ scores are familiar to many practitioners and have a 1-1 correspondence with $p$ values. The plotted $Z$ scores are the number of standard deviations from the mean of a standard normal such that the left and right tails together have probability $p$. For comparison, a test with a single decision point that spends all of $\alpha=0.05$ on that decision point would require $Z\approx 1.96$ for $p=0.05$.) The $Z$ score required for significance for that method is
@@
with $\rho$ a parameter that, like $w$ for our method, controls the $t$ value at which the $Z$ score reaches a minimum. Both methods apply to unbounded tests.
@@
Figure 2: Comparison to another always-valid method. For $u=0.2$, requiring repetition gives less stringent requirements for significance than a method based on autocorrelation for averages waudbysmith24 .
5 Nearly Impossible is Possible
Figure 2 has $Z$ score curves that increase after reaching a minimum, prompting the question: “Is there a way to spend $\alpha$ that gives a flat curve for an unbounded test, keeping the required $p$-value from shrinking to zero, or, equivalently, the required $Z$ score from becoming unbounded as $t\rightarrow\infty$?"
@@ 5 Nearly Impossible is Possible
Now we show that the answer to the question is a negative of the form: “It is not quite possible, but nearly so."
@@
$\frac{c}{\alpha u}\left(\frac{1}{t+\frac{1}{u}}\right).$ (35)
@@
diverges for $0\leq v\leq 1$ and converges for $v>1$. So in Expression 38, the first sum diverges and the second converges, implying that the expression diverges. ∎
In short, we cannot have $x_{t}$ proportional to $\frac{1}{t}$ because its $p$-series diverges. However, for any $v>1$, we can use $x_{t}$ proportional to $\frac{1}{t^{v}}$, since its $p$-series converges. So, in a sense, Theorem 3 is the most nearly possible impossible result possible.
@@
$x_{t}=\frac{1}{\left[\zeta(v)-h(s,v)\right]\left(t+s\right)^{v}},$ (46)
so
$\delta_{t}=\frac{1}{\left[\zeta(v)-h(s,v)\right]\left(t+s\right)^{v}}\alpha\lceil ut\rceil.$ (47)
Figure 3 shows the results of removing some initial terms from the $p$-series to produce $x_{t}$. The curve for $s=0$ has no terms removed. Note that removing more terms lowers the required $Z$ scores for significance for most of the $t$ values.
@@
6 Discussion
We have explored different ways to spend $\alpha$ with repetition requirements for unbounded tests, showing that it is possible to adjust the levels of significance required at each decision point by requiring different amounts of repetition. We have also shown that it is impossible to have a significance requirement that does not eventually become more stringent for an unbounded test with a repetition requirement that is a fixed fraction of the prior decision points, but we can stave that off for an arbitrary number of decision points by setting parameters appropriately.
An earlier paper bax_sarkar_shtoff that focused on bounded tests discussed how to balance control of Type I and Type II error, by spending a portion of $\alpha$ on a single loose requirement for significance at the end of the test, which is reminiscent of traditional sequential testing strategies, such as the Haybittle-Peto boundary haybittle71 ; peto76 . It is a challenge to imagine something similar for unbounded tests, since their ending decision points are not known a priori. Perhaps spending a portion of $\alpha$ for loose requirements at a (possibly unbounded) subsequence of the decision points, either with or without a repetition requirement, would prove useful.
It would also be interesting to explore ways to combine repetition requirements with other approaches to continuous monitoring and always-valid bounds deng16 ; johari21 ; maharaj23 ; grunwald24 ; waudbysmith24 . In some cases, it may be possible to improve both types of bounds by combining their strategies. The challenge is that non-repetition methods sometimes make use of side conditions that are instead proven empirically in some sense with repetition requirements – we need to analyze how these approaches overlap to understand how they can complement each other beyond mere duplication, if possible.